OpenStudy (anonymous):

System of Equations, solve for x & y 3x+4y=4 6x+4y=16

OpenStudy (anonymous):

@ganeshie8 @billj5 @Mashy @NaomiBell1997 @Goku-Kai @King.Void. @just_one_last_goodbye

OpenStudy (anonymous):

ok so solve one of the 2 for x or y, you pick, since you wanted to use substitution

OpenStudy (anonymous):

subtract 1st eq. from 2nd and find the value of x and then y

OpenStudy (anonymous):

one of the two equations i mean

OpenStudy (anonymous):

yeah elimination might be easier for this one

OpenStudy (anonymous):

Elimination is the easiest. Do you know how? If not, you subtract the equations. So if you have 3x + 2 = 20 and then you have 1x + 1 = 10 the answer would be 2x + 1 = 10/ Make sense? Please tell me if it doesn't. :)

OpenStudy (anonymous):

Oops, forgot to finish how to do it lol. so you have 2x + 1 = 10 just do as you would normally do. 2x + 1 - 1 = 10 - 1 2x = 9 2x/2 = 9/2 x=4.5 Now substitute that into the equation 3x + 2 = 20 3(4.5) + 2 = 20 see? (I know the problem doesn't equal 20=20 but you get the point lol) :)

OpenStudy (anonymous):